# Temporally Compressed 3D Gaussian Splatting for Dynamic Scenes

### v1

@@ 9.2 Quantization
9.3 Keypoint Interpolation
@@
Dynamic 3DGS [26] Pruned Quantized + Pruned Quantized + Pruned + KPI
@@
Table 8: Per-scene quantitative comparisons on the Neural 3D Video Dataset [22]. Some methods only report part of the scenes. 1 only includes the Flame Salmon scene. 2 excludes the Coffee Martini scene. “-” denotes results that are unavailable in prior work.
[…]
Neural Volumes [25] 1 $22.80$ $22.80$
LLFF [27] 1 $23.24$ $23.24$
DyNeRF [22] 1 $29.58$ $29.58$
HexPlane [3] 2 $31.71$ $32.04$ $32.55$ $29.47$ $32.08$ $32.39$
NeRFPlayer [36] $30.69$ 31.53 $30.56$ $29.35$ $31.65$ $31.93$ $29.13$
HyperReel [1] $31.10$ $28.37$ $32.30$ $32.92$ $28.26$ $32.20$ $32.57$
K-Planes [16] $31.63$ $29.99$ $32.60$ $31.82$ $30.44$ $32.38$ $32.52$
MixVoxels-L [37] $31.34$ $29.63$ $32.25$ $32.40$ $29.81$ $31.83$ $32.10$
MixVoxels-X [37] $31.73$ $30.39$ $32.31$ $32.63$ 30.60 $32.10$ $32.33$
4D Gaussians [38] $30.67$ $27.34$ $32.46$ $32.90$ $29.20$ $32.51$ $32.49$
Dynamic 3DGS [26] $30.97$ $27.32$ $32.97$ $31.75$ $27.26$ $33.24$ $33.68$
STG [23] 32.05 $28.61$ 33.18 $33.52$ $29.48$ 33.64 33.89
Ours $30.58$ $27.42$ $31.92$ 33.61 $27.12$ $32.48$ $33.04$
[…]
Neural Volumes [25] 1 $0.295$ $0.295$
LLFF [27] 1 $0.235$ $0.235$
DyNeRF [22] 1 $0.083$ $0.083$
HexPlane [3] 2 $0.075$ $0.082$ $0.080$ $0.078$ $0.066$ $0.070$
NeRFPlayer [36] $0.111$ $0.085$ $0.113$ $0.144$ $0.098$ $0.088$ $0.138$
HyperReel [1] $0.096$ $0.127$ $0.089$ $0.084$ $0.136$ $0.078$ $0.077$
MixVoxels-L [37] $0.096$ $0.106$ $0.099$ $0.088$ $0.116$ $0.088$ $0.080$
MixVoxels-X [37] $0.064$ $0.081$ $0.062$ $0.057$ $0.078$ $0.051$ $0.053$
Dynamic 3DGS [26] $0.082$ $0.122$ $0.077$ $0.087$ $0.09$ $0.072$ $0.068$
STG [23] 0.044 0.069 0.037 0.036 0.063 0.029 0.030
Ours $0.089$ $0.129$ $0.081$ $0.088$ $0.094$ $0.075$ $0.071$
@@
Figure 12: Qualitative Evaluation on the Neural 3D Video Dataset. This figure shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near-identical results despite extreme compression. Top: Cook Spinach, Bottom: Sear Steak

### v2

@@ 5.3 Mixed-Precision Quantization
6 Conclusion
We have introduced Temporal Compressed 3D Gaussian Splatting (TC3DGS), a novel framework designed to achieve memory-efficient and high-speed reconstruction of dynamic scenes. Our approach achieves up to 67x compression and up to three times faster rendering while maintaining high fidelity in complex motion scenarios. Through selective temporal pruning and gradient-based quantization, TC3DGS significantly reduces memory usage with minimal impact on visual quality. While our method is effective, limitations remain in aggressively compressing position data and handling new objects that enter the scene mid-sequence. Future work will focus on bridging the gap between spatio-temporal methods and storage-efficient dynamic scene reconstruction, improving adaptability and further reducing storage requirements for complex, evolving environments.
Supplementary
7 Implementation Details
For our experiments, we closely follow the hyperparameters outlined in Dynamic 3DGS [luiten2023dynamic]. Specifically, for the compression strategy, we set both the mask weight parameter $\lambda_{\text{mask}}$ and the masking consistency loss parameter $\lambda_{\text{mask-cons}}$ to $0.01$ in our primary experiments. In terms of quantization, we use Gaussian parameters with bin sizes of $b_{\text{min}}=4$ and $b_{\text{max}}=8$, while positional data is quantized to 16-bit precision to ensure spatial accuracy. For the learnable quantization step size, we use a $0.02$ learning rate. Quantization is applied after 6000 iterations in the first scene for all experiments. For keypoint interpolation, we set the tolerance to $\tau=1\times 10^{-5}$ for the Panoptic dataset and $\tau=1\times 10^{-7}$ otherwise. Moreover, we restrict the maximum number of keypoints to $\text{max}_{\text{kp}}=30$ for the Panoptic dataset and $\text{max}_{\text{kp}}=60$ otherwise, which provides an effective balance between compression and temporal trajectory accuracy. Additional ablation studies with varying values of hyperparameters are provided in the supplementary.
8 Additional Ablation Experiments
We conducted several experiments to better understand hyperparameters of our method as shown in Figure 6.
@@ 8 Additional Ablation Experiments
Dynamic 3DGS [luiten2023dynamic] Pruned Quantized + Pruned Quantized + Pruned + KPI
@@
8.1 Masking and Mask Consistency
(a) SSIM $\uparrow$
Mask Consistency Loss
Loss 0 0.01 0.05 0.1
0.005 0.909 0.911 0.911 0.907
0.01 0.913 0.914 0.911 0.909
0.05 0.904 0.901 0.899 0.897
0.1 0.894 0.896 0.895 0.897
(b) PSNR $\uparrow$
Mask Consistency Loss
Loss 0 0.01 0.05 0.1
0.005 28.0 27.8 28.2 26.8
0.01 28.1 28.3 26.8 26.9
0.05 27.2 26.6 26.0 26.0
0.1 27.1 27.4 26.6 26.7
(c) LPIPS $\downarrow$
Mask Consistency Loss
Loss 0 0.01 0.05 0.1
0.005 0.180 0.177 0.178 0.182
0.01 0.174 0.176 0.175 0.183
0.05 0.205 0.209 0.214 0.217
0.1 0.235 0.229 0.234 0.234
(d) Number of Gaussians $\downarrow$
Mask Consistency Loss
Loss 0 0.01 0.05 0.1
0.005 208k 156k 143k 141k
0.01 154k 113k 98k 94k
0.05 50k 47k 30k 26k
0.1 29k 31k 19k 16k
@@ 8.1 Masking and Mask Consistency
8.2 Keypoint Interpolation
9 Detailed Results for Panoptic and Neural 3D Video Dataset
We present detailed results for the two datasets discussed in Section 5.1 of our paper, summarized in Table 5 and 6. These results include various metrics evaluated for each scene in both datasets, offering a granular view of performance. Additionally, we provide supplementary visualizations for the Neural 3D Video dataset, which further demonstrate the effectiveness of our method in capturing dynamic scenes with high fidelity. Finally, we present an ablation study over the main elements of our method in Figure˜6.
Table 5: Per-scene quantitative comparisons on the Panoptic Sports Dataset [Joo_2019].
Method Avg. Basket. Juggle Boxes Softball Tennis Football
[…]
4DGaussians [4dgs] $27.2$ $26.68$ $27.47$ $27.02$ $27.44$ $27.31$ $27.39$
STG [stg] $20.45$ $21.60$ $19.93$ $20.65$ $19.44$ $20.82$ $20.23$
D. 3DGS [luiten2023dynamic] 28.70 28.22 29.48 29.46 28.43 28.11 $28.49$
HiCom [hicom2024] $26.59$ $26.35$ $27.15$ $25.09$ $28.38$ $25.91$ 28.62
Ours $27.81$ $28.02$ $28.65$ $28.28$ $27.96$ $25.97$ $28.00$
[…]
4DGaussians [4dgs] 0.91 $0.90$ 0.92 0.91 0.92 0.92 0.92
STG [stg] $0.79$ $0.78$ $0.78$ $0.79$ $0.79$ $0.78$ $0.78$
D. 3DGS [luiten2023dynamic] 0.91 0.91 0.92 0.91 $0.91$ $0.91$ $0.91$
HiCoM [hicom2024] $0.87$ $0.82$ $0.88$ $0.86$ $0.90$ $0.86$ $0.91$
Ours $0.89$ $0.89$ $0.90$ $0.89$ $0.89$ $0.89$ $0.89$
[…]
4DGaussians [4dgs] $63$ $66$ $59$ $63$ $57$ $72$ $65$
STG [stg] 19 19 19 20 19 22 19
D. 3DGS [luiten2023dynamic] $2008$ $2161$ $1935$ $2021$ $2021$ $1915$ $2000$
HiCoM [hicom2024] $71$ $71$ $71$ $76$ $70$ $69$ $70$
Ours $49$ $46$ $32$ $44$ $41$ $48$ $34$
Table 6: Per-scene quantitative comparisons on the Neural 3D Video Dataset [li2022neural3dvideosynthesis]. Some methods only report part of the scenes. 1 only includes the Flame Salmon scene. 2 excludes the Coffee Martini scene. “-" denotes results that are unavailable in prior work. STG on the Neural 3D Video dataset is trained on 50-frame sequences and requires six models for evaluation.
[…]
Neural Volumes [Lombardi2019] 1 $22.80$ - - - $22.80$ - -
LLFF [mildenhall2019llff] 1 $23.24$ - - - $23.24$ - -
DyNeRF [li2022neural3dvideosynthesis] 1 $29.58$ - - - $29.58$ - -
HexPlane [Cao2023HexPlane] 2 $31.71$ - $32.04$ $32.55$ $29.47$ $32.08$ $32.39$
NeRFPlayer [song2023nerfplayer] $30.69$ 31.53 $30.56$ $29.35$ $31.65$ $31.93$ $29.13$
HyperReel [attal2023hyperreel] $31.10$ $28.37$ $32.30$ $32.92$ $28.26$ $32.20$ $32.57$
K-Planes [kplanes_2023] $31.63$ $29.99$ $32.60$ $31.82$ $30.44$ $32.38$ $32.52$
MixVoxels-L [Wang2023ICCV] $31.34$ $29.63$ $32.25$ $32.40$ $29.81$ $31.83$ $32.10$
MixVoxels-X [Wang2023ICCV] $31.73$ $30.39$ $32.31$ $32.63$ 30.60 $32.10$ $32.33$
4D Gaussians [4dgs] $30.67$ $27.34$ $32.46$ $32.90$ $29.20$ $32.51$ $32.49$
Dynamic 3DGS [luiten2023dynamic] $30.97$ $27.32$ $32.97$ $31.75$ $27.26$ $33.24$ $33.68$
STG [stg] 32.05 $28.61$ 33.18 $33.52$ $29.48$ 33.64 33.89
Ours $30.96$ $27.42$ $32.72$ 33.61 $27.12$ $32.48$ $33.04$
[…]
Neural Volumes [Lombardi2019] 1 $0.295$ - - - $0.295$ - -
LLFF [mildenhall2019llff] 1 $0.235$ - - - $0.235$ - -
DyNeRF [li2022neural3dvideosynthesis] 1 $0.083$ - - - $0.083$ - -
HexPlane [Cao2023HexPlane] 2 $0.075$ - $0.082$ $0.080$ $0.078$ $0.066$ $0.070$
NeRFPlayer [song2023nerfplayer] $0.111$ $0.085$ $0.113$ $0.144$ $0.098$ $0.088$ $0.138$
HyperReel [attal2023hyperreel] $0.096$ $0.127$ $0.089$ $0.084$ $0.136$ $0.078$ $0.077$
MixVoxels-L [Wang2023ICCV] $0.096$ $0.106$ $0.099$ $0.088$ $0.116$ $0.088$ $0.080$
MixVoxels-X [Wang2023ICCV] $0.064$ $0.081$ $0.062$ $0.057$ $0.078$ $0.051$ $0.053$
Dynamic 3DGS [luiten2023dynamic] $0.082$ $0.122$ $0.077$ $0.087$ $0.09$ $0.072$ $0.068$
STG [stg] 0.044 0.069 0.037 0.036 0.063 0.029 0.030
Ours $0.089$ $0.129$ $0.081$ $0.088$ $0.094$ $0.075$ $0.071$
10 Technicolor Dataset [technicolor]
To further demonstrate the effectiveness of our method, we evaluate our method on the Technicolor Dataset. The results in Table 7 show that our method provides competitive performance with a significantly smaller storage footprint. Figure˜9 shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near identical results.
[…]
Method Avg. Birthday Fabien Painter Theater Train
[…]
STG [stg] $33.60$ $32.09$ $35.70$ $36.44$ $30.99$ $32.58$
D. 3DGS [luiten2023dynamic] $32.12$ $30.82$ $33.62$ $34.73$ $31.12$ $30.67$
Ours $31.9$ $30.33$ $33.15$ $34.62$ $31.01$ $30.57$
[…]
STG [stg] - - - - - -
D. 3DGS [luiten2023dynamic] $0.88$ $0.90$ $0.88$ $0.87$ $0.86$ $0.88$
Ours $0.85$ $0.88$ $0.86$ $0.81$ $0.84$ $0.86$
[…]
STG [stg] $0.08$ $0.04$ $0.11$ $0.10$ $0.13$ $0.04$
D. 3DGS [luiten2023dynamic] $0.13$ $0.07$ $0.16$ $0.16$ $0.14$ $0.11$
Ours $0.14$ $0.09$ $0.15$ $0.17$ $0.16$ $0.13$
11 Limitations
@@
Figure 10: Qualitative Evaluation on the Neural 3D Video dataset. This figure shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near-identical results despite extreme compression. Top: Cook Spinach, Bottom: Sear Steak.
11.1 Datasets
11.1.1 Panoptic Sports Dataset
We evaluate our method on the Panoptic Sports dataset, a subset of the Panoptic Studio dataset [Joo_2019]. This dataset contains 6 different scenes, each having 31 camera sequences spanning 150 frames. We use 4 cameras (0, 10, 15 and 30) for testing and the rest for training, following the convention set by [luiten2023dynamic]. In addition to the images, we use the provided foreground/background masks to apply a segmentation loss to improve the results and prevent the background from moving. This dataset contains complex motions with objects moving quickly and over long distances.
11.1.2 Neural 3D Video Dataset
[…]